Solve This system of equations:
2x-3x=13
x+y=-1

Answer by Ave(106)

2x-3x=13
x+y=-1
I'll be using substitution
2x-3x=13
-x=13
x=-13
Plug in -13 for x in the other equation to find y
-13+y=-1
y=12
(-13, 12)